At approximately 10:45 AM, Friday, November 7, 2008, new Internet requests from the CoC began failing. This affected CoC faculty, staff, and students. Working with OIT, it was determined that a recently-made firewall configuration change was not freeing up connections behind the North Interconnect router quickly enough. So, while existing connections continued to function, new connections were being refused. OIT reversed the change, and services were returned to normal at 11:30 AM, November 7.